OpenStudy (anonymous):

PLEASE HELP !! which ratio forms a proportion with 24/20 ? A. 6/4 B. 10/12 C. 6/5 D. 3/5

OpenStudy (aaronq):

what is the largest number you can divide 24/20 by that will give you integers?

OpenStudy (calculusxy):

Use the GCF method.

OpenStudy (anonymous):

whats gcf method ?

OpenStudy (anonymous):

where num is a constant ?*num=24 ?*num=20 gcf is greatest common factor

OpenStudy (calculusxy):

Since the GCF is 4, we know that if we divide 24÷4=6 and 20÷4=5. So the proportion will be 6/5

OpenStudy (anonymous):

so the answer is b ?

OpenStudy (anonymous):

No It's C

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

Since the GCF is 4, we know that if we divide 24÷4=6 and 20÷4=5. So the proportion will be 6/5

OpenStudy (anonymous):

ok

OpenStudy (calculusxy):

The GCF method is the greatest common factor method.